OpenStudy (anonymous):

Use the Substitution Method to solve the following system of equations. 2x + y = 3 -3x + 4y = 1 (1, 1) (-1, -1) (1, -1) (-1, 1)

OpenStudy (itsonlycdeee):

2x+y =3 becomes y = -2x +3 so you plug that into -3x+4y=1 --> -3x + 4(-2x)+3. Can you do the rest?

OpenStudy (anonymous):

-3 so 1(-2x)?

OpenStudy (anonymous):

x=1, y=1

OpenStudy (anonymous):

^ thats the answer?

OpenStudy (itsonlycdeee):

Sorry I meant -3x + 4(-2x+3) = 1

OpenStudy (anonymous):

but is that the answer?

OpenStudy (itsonlycdeee):

Yes

OpenStudy (anonymous):

thanks

OpenStudy (itsonlycdeee):

No problem!